OpenStudy (anonymous):

solve the system by elimination method..thanks guys 4x=18+2y -6/5x+y=11/5

sam (.sam.):

4x-2y=18 -6/5x+y=11/5......multiply-2 --------------- 4x-2y=18 12x/5-2y=-22/5 8x/5=112/5 8x=112 x=14 ------------ 4x-2y=18 4(14)-2y=18 56-2y=18 -2y=-38 y=19

OpenStudy (anonymous):

wowo sam you are my hero..<3